OpenStudy (anonymous):

using the formulas for lowering powers to rewrite the expression in terms of first power of cosine cos(x)^7 ..can anyone help me with this?

OpenStudy (anonymous):

Cos^2(x) = (1+Cos(2x)) /2 Cos^7(x) = Cos(x) * [(1+Cos(2x)) /2] ^3 then expand and further simplify terms

OpenStudy (anonymous):

how do you expand it?